# PP-Tac: Paper Picking Using Tactile Feedback in Dexterous Robotic Hands

### v1

@@ V-A Grasp Motion Dataset Synthesis
We synthesize grasping motions through trajectory optimization in simulation, avoiding the need for complex teleoperation devices. While reinforcement learning (RL) offers an alternative, it requires soft-body simulation to model deformable object dynamics and VBTS elastomer behavior, often necessitating additional real-to-sim procedures for fidelity. In contrast, our approach uses rigid-body dynamics and transfers directly to real robots, as validated experimentally. The grasping process begins by establishing contact between the fingertips and the object’s surface (see Section -B for implementation details). Once contact is achieved, the fingers gradually close to complete the grasp. Each finger follows an independent trajectory on the object’s surface, with target normal forces(Figs. 4 and 5).
@@
To account for varying material properties, sliding trajectories are adjusted based on fingertip contact forces $F$. This is achieved by synthesizing motions that deform the tactile sensor’s elastomer layer to different extents. The deformation reading $\boldsymbol{d}_{tac}$ is proportional to the applied pressure, governed by the elastomers Young’s modulus. Thus, contact force $F$ is modulated by controlling $\boldsymbol{d}_{tac}$ via position control. Notably, the exact relationship between $\boldsymbol{d}_{tac}$ and $F$ is not explicitly modeled, as precise force values are unnecessary for the algorithm. Our approach leverages rigid-body dynamics to control contact forces efficiently, avoiding complex deformable dynamics calculations. By adjusting the distance between the finger joint and the terrain, we can obtain trajectories with varying degrees of deformation. When the distance between the finger joint and the terrain is equal to the sensor’s radius as shown Fig. 5, the finger just makes contact with the terrain and $\boldsymbol{d}_{tac}$ just equals to 0. Finally, we get the target trajectories of 4 fingertips with variant contact conditions denoted as $\boldsymbol{ee}_{target}$
@@ V-B PP-Tac Policy
Once the dataset is prepared, we employ a diffusion policy to jointly control the hand and arm, enabling adaptation to varying terrain shapes and contact force conditions. We adopt a Denoising Diffusion Probabilistic Model (DDPM) framework [17, 18, 8, 41], which predicts future actions ($N_{pred}$ steps of $x^{pred}$) conditioned on historical states ($N_{prefix}$ steps of $x^{prefix}$). In each frame, the state variables include:
@@
where $\boldsymbol{p}\in\mathbb{R}^{17\times 3}$ is hand joints’ position in world coordinate, $\dot{\boldsymbol{p}}\in\mathbb{R}^{17\times 3}$ is the linear velocity of the hand joints relative to each parent frame, $\boldsymbol{q}\in\mathbb{R}^{16}$ is the rotation angle of controllable hand joints, $\boldsymbol{\dot{q}}\in\mathbb{R}^{16}$ is the angular velocity of controllable hand joints, $R\in\mathbb{R}^{6}$ is 6D rotation (represented as two-row vectors of rotational martix, which is from [50]) of wrist(end effector of arm), $\Omega\in\mathbb{R}^{6}$ represents the angular velocity of wrist rotation, $p_{wrist}\in\mathbb{R}$ is the wrist’s height along arm’s $z$-axis, $\dot{p}_{wrist}\in\mathbb{R}$ is the linear velocity of $p_{wrist}$, $\boldsymbol{d}_{tac}\in\mathbb{R}^{4}$ represents the deformation depth readings from four fingertip tactile sensors. Table II summarizes the notations used in this paper. The total state dimension is $\mathcal{D}=152$. Such a over-parameterized input allows the network to extract more robust and expressive latent features for the diffusion policy.
The pipeline is illustrated in Fig. 6 and Fig. 6 (right) illustrates a single denoising diffusion step. We apply an encoder-only transformer to predict future robot motion $x_{0}^{pred}$ given prefix motion $x^{prefix}$, diffused future motion $x_{t}^{pred}$, diffusion step $t$, current frame index $i$, and target deformation depth $\boldsymbol{\bar{d}}_{tac}$. The input sequence is encoded into a latent vector of dimension $\mathbb{R}^{(1+N_{prefix}+N_{pred})\times\mathcal{D}}$, comprising: 1) A latent vector of $\mathcal{D}$-dimensional features representing $t$, $i$, and $\boldsymbol{\bar{d}}_{tac}$, extracted using three 3-layer MLP networks respectively. 2) $N_{prefix}\times\mathcal{D}$ dimensions corresponding to the prefix states of $N_{prefix}$ time steps. 3) $N_{pred}\times\mathcal{D}$ dimensions for the predicted states of $N_{pred}$ time steps. Instead of predicting $\epsilon_{t}$ as formulated by [18], we follow [45] to predict the state sequence itself $\hat{x}_{0}^{pred}$. Predicting $\hat{x}_{0}^{pred}$ is found to produce better results for the state sequence which contains motion data, and enables us to apply a target loss as geometric loss explicitly as each denoising step as following:
@@ VI-D Comparison with Other System Configurations
To assess whether PP-Tac’s system setup leveraging dexterous hand and tactile sensors can offer advantages, systematic comparisons with other robot configurations were conducted. Here, we constructed three baselines:
@@
To ensure fairness, each trial allowed only one grasp attempt.
@@ IX Acknowledgment
We extend our thanks to Changyi Lin (Carnegie Mellon University) for insightful discussions.

### v2

@@ V-A Grasp Motion Dataset Synthesis
We synthesize grasping motions through trajectory optimization in simulation, avoiding the need for complex teleoperation devices. While reinforcement learning (RL) offers an alternative, it requires soft-body simulation to model deformable object dynamics and VBTS elastomer behavior, often necessitating additional real-to-sim procedures for fidelity. In contrast, our approach uses rigid-body dynamics and transfers directly to real robots, as validated experimentally. The grasping process begins by establishing contact between the fingertips and the object’s surface (see Section -B for implementation details). Once contact is achieved, the fingers gradually close to pinch the object. Each finger follows an independent trajectory on the object’s surface, and concurrently exerts target normal forces (Figs. 4 and 5).
@@
To account for varying material properties, sliding trajectories are adjusted based on fingertip contact forces $F$. This is achieved by synthesizing motions that deform the tactile sensor’s elastomer layer to different extents. The deformation reading $\boldsymbol{d}_{tac}$ is proportional to the applied pressure, governed by the elastomers Young’s modulus. Thus, contact force $F$ is modulated by controlling $\boldsymbol{d}_{tac}$ via position control. Notably, the exact relationship between $\boldsymbol{d}_{tac}$ and $F$ is not explicitly modeled, as precise force values are unnecessary for the algorithm. Our approach leverages rigid-body dynamics to control contact forces efficiently, avoiding complex deformable dynamics calculations. By adjusting the distance between the finger joint and the terrain, we can obtain trajectories with varying degrees of deformation. For example, when the distance between the finger joint and the terrain is equal to the sensor’s radius (Fig. 5), the finger just makes contact with the terrain and $\boldsymbol{d}_{tac}$ just equals to 0. Finally, we get the target trajectories of 4 fingertips, with their positions denoted as $\boldsymbol{ee}_{target}$.
@@ V-B PP-Tac Policy
Once the dataset is prepared, we employ a diffusion policy to jointly control the hand and arm, enabling adaptation to varying terrain shapes and contact force conditions. We adopt a Denoising Diffusion Probabilistic Model (DDPM) framework [DDPM, 16, 8, 39], which predicts future actions ($N_{pred}$ steps of $x^{pred}$) conditioned on historical states ($N_{prefix}$ steps of $x^{prefix}$). In each frame, the state variables include:
@@
where $\boldsymbol{p}\in\mathbb{R}^{17\times 3}$ is hand joints’ position in world coordinate, $\dot{\boldsymbol{p}}\in\mathbb{R}^{17\times 3}$ is the linear velocity of the hand joints relative to each parent frame, $\boldsymbol{q}\in\mathbb{R}^{16}$ is the rotation angle of controllable hand joints, $\boldsymbol{\dot{q}}\in\mathbb{R}^{16}$ is the angular velocity of controllable hand joints, $R\in\mathbb{R}^{6}$ is 6D rotation (represented as two-row vectors of rotational martix, which is from [Zhou_2019_CVPR]) of wrist(end effector of arm), $\Omega\in\mathbb{R}^{6}$ represents the angular velocity of wrist rotation, $p_{wrist}\in\mathbb{R}$ is the wrist’s height along arm’s $z$-axis, $\dot{p}_{wrist}\in\mathbb{R}$ is the linear velocity of $p_{wrist}$, $\boldsymbol{d}_{tac}\in\mathbb{R}^{4}$ represents the deformation depth readings from four fingertip tactile sensors. Table II summarizes the notations used in this paper. The total state dimension is $\mathcal{D}=152$. Such a over-parameterized input allows the network to extract more robust and expressive latent features for the diffusion policy.
The pipeline is illustrated in Fig. 6. Fig. 6 (right) illustrates a single denoising diffusion step. We apply an encoder-only transformer to predict future robot motion $x_{0}^{pred}$ given prefix motion $x^{prefix}$, diffused future motion $x_{t}^{pred}$, diffusion step $t$, current frame index $i$, and target deformation depth $\boldsymbol{\bar{d}}_{tac}$. The input sequence is encoded into a latent vector of dimension $\mathbb{R}^{(1+N_{prefix}+N_{pred})\times\mathcal{D}}$, comprising: 1) A latent vector of $\mathcal{D}$-dimensional features representing $t$, $i$, and $\boldsymbol{\bar{d}}_{tac}$ extracted using a three 3-layer MLP network. 2) $N_{prefix}\times\mathcal{D}$ dimensions corresponding to the prefix states of $N_{prefix}$ time steps. 3) $N_{pred}\times\mathcal{D}$ dimensions for the predicted states of $N_{pred}$ time steps. Instead of predicting $\epsilon_{t}$ (formulated by [16]), we follow [43] to predict the state sequence itself $\hat{x}_{0}^{pred}$. Predicting $\hat{x}_{0}^{pred}$ is found to produce better results for the state sequence which contains motion data, and enables us to apply a target loss for each denoising step as following:
@@ VI-D Comparison with Other System Configurations
To assess whether PP-Tac’s system setup leveraging dexterous hand and tactile sensors can offer advantages, systematic comparisons with other robot configurations were conducted. Here, we constructed three baselines. To ensure fairness, each trial allowed only one grasp attempt.
@@ IX Acknowledgment
We thank Changyi Lin (Carnegie Mellon University) for insightful discussions. This work was supported in part by the National Natural Science Foundation of China (Grant No.52305007), by the State Key Laboratory of Mechanical System and Vibration (Grant No. MSV202519), and by Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI), MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (KLIP-HuMaCo).